When operated on a household 110.0-V line, typical hair dryers draw about 1550 W of power. We can model the current as a long straight wire in the handle. During use, the current is about 2.25 cm from the user\'s hand. What is the current in the dryer? What is the resistance of the dryer? What magnetic field does the dryer produce at the user\'s hand?
Answer #1

Given:

V = 110V

P = 1550W

r = 2.25cm

We know, P = V2/R

=> R = V2/P

= (110)2/1550

= 7.806 ohm

B = μ0I/2πr

First, I = P/V = 1550/110 = 14.09A

So,

B = (4π*10-7)(14.09)/(2π*0.0225)

= 1.25 x 10-4 T